# QVGen: Pushing the Limit of Quantized Video Generative Models

### v4

@@ Appendix F Additional Fine-Grained Decay Strategies
In this section, we detail the “Sparse” and “Res. Q.” decay strategies (see Tab. R), which shrink $\Phi$ to $\varnothing$, as follows:
@@
    $\begin{array}[]{ll}{\mathbf{W}}_{\Phi}={\mathcal{Q}}_{4}({\mathbf{W}}_{\Phi})+{\mathcal{Q}}_{4}({\mathbf{E}}_{1})+{\mathcal{Q}}_{4}({\mathbf{E}}_{2})+{\mathcal{Q}}_{4}({\mathbf{E}}_{3}).\end{array}$ (I)
@@
Compared with our rank-decay strategy (with $r=32$), both alternative methods require storing multiple tensors [^8] The “Sparse” strategy additionally requires storing binary masks to enforce sparsity., each with the same shape as the corresponding linear layer’s weight matrix. In particular, the “Res. Q.” strategy incurs substantial memory overhead and necessitates CPU offloading to avoid out-of-memory (OOM) issues. Furthermore, the “Sparse” strategy introduces extra computational cost due to the matrix multiplication between the mask and ${\mathbf{W}}_{\Phi}$. In practice, the “Sparse” and “Res. Q.” strategies consume $20.12$ and $28.78$ GPU days, respectively, whereas our proposed rank-decay requires only $11.11$ GPU days (see Tab. G). More importantly, rank-decay also achieves significantly better performance than both alternatives, as demonstrated in Tab. R.
Appendix G Further Analyses of Gradient Norm in Video Generation QAT
G.1 Video Generation QAT vs. Image Generation QAT
Table I: Average gradient norm comparison. SD3-medium (Esser et al., 2024a) is an advanced diffusion model for image generation. We train SD3-medium for $2K$ steps with $16K$ images from the LAION-55B dataset (Schuhmann et al., 2022) on 88 $H100$ GPUs. “Avg. $\|{\bm{g}}_{t}\|_{2}$” is the average of the gradient norm across training steps.
@@ Appendix G Further Analyses of Gradient Norm in Video Generation QAT
First, we conduct experiments to compare the gradient norm between image generation QAT and video generation QAT. In Tab. I, we employ W4A4 Q-DM (Li et al., 2023b) to quantize the diffusion models under the same settings. We observe that reducing the gradient norm during QAT, ignored by previous research (Li et al., 2023b; He et al., 2024; Esser et al., 2020b), is far more critical for video generation QAT than for image generation QAT. As Tab. I shows, with a similar parameter count and the same QAT recipe, video diffusion reaches a significantly larger gradient norm than image diffusion. This leads to much more unstable optimization (Xie et al., 2024) in training. We believe the phenomenon happens because video generation introduces complicated temporal modeling, which eventually makes quantization for video diffusion more challenging than image diffusion.
G.2 Impact of Motion Dynamics on Gradient Norm
Table J: Results between different training videos. We employ W4A4 QVGen with the same configurations as those in Tab. 1 to quantize CogVideoX-22(Yang et al., 2025).
@@ G.2 Impact of Motion Dynamics on Gradient Norm
Here, we study how the motion dynamics of video generation affect the gradient norm. Specifically, using UniMatch (Xu et al., 2023a), we compute an optical‑flow score as a motion difference score for each clip and split the training data (Sec. 4.1) into high‑motion and low‑motion subsets, each with $8K$ videos. In Tab. J, the model trained on the low‑motion subset shows a much lower average gradient norm and a better Imaging Quality score, but its Dynamic Degree is lower than that of the model trained on the high‑motion subset. This pattern suggests that a high-motion training set makes QAT harder and less stable (i.e., higher gradient norm), lowering static quality but boosting motion quality. The reverse holds for low‑motion clips. To be noted, with a mixed dataset (containing low-motion ++ high-motion) in the $4$-th row of Tab. J, the quantized model generalizes well in producing either high‑motion or low‑motion content. Therefore, although a low-motion dataset in training can cause slightly lower $\|{\bm{g}}_{t}\|_{2}$ (i.e., better training convergence), it is necessary to properly add high-motion data to improve motion quality.
Appendix H Comparison with baselines on Additional Metrics
Table K: Additional W4A4 performance comparison across different quantization methods. We employ the same models as those in Tab. 1.
@@ Appendix H Comparison with baselines on Additional Metrics
We also evaluate the similarity between videos generated by different quantization methods and those generated by BF16 models on VBench (Huang et al., 2024b) captions. Specifically, we employ PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity) (Wang et al., 2004), and LPIPS (Learned Perceptual Image Patch Similarity) (Zhang et al., 2018). In Tab. K, QVGen substantially outperforms the baselines on these metrics.
Table L: Performance comparison for huge DMs across different methods on VBench (Huang et al., 2024b). We employ W4A4 CogVideo-X 55(Yang et al., 2025) here.
Method
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Appendix I Comparison with Baselines for Huge DMs
We include a comparison on VBench for large-scale CogVideoX‑1.5 55B in Tab. L. QVGen again surpasses all baselines, just as it does on smaller models, underscoring its strength across a wide range of model sizes. Limited resources currently prevent us from adding more baselines for these large-scale models, but we plan to do so in future work.
Appendix J More Ablation Studies
J.1 Robustness of Rank-Decay Schedule across Different Annealing Functions
In this section, we test $5$ different annealing functions for $u$. The results in Tab. M reveal that all the functions yield comparable performance, highlighting the robustness of our approach.
Table M: Results of different annealing factors $u$. All of them decay from $1$ to 0. We employ “Cosine” in this work.
$\quad u\quad$
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@ Appendix I Comparison with Baselines for Huge DMs
J.2 Initialization Methods for Auxiliary Modules $\Phi$
Table N: W4A4 results of different initialization strategies for ${\mathbf{W}}_{\Phi}$.
Init.
Strategy
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Besides employing ${\mathbf{W}}-{\mathcal{Q}}_{b}({\mathbf{W}})$ to initialize ${\mathbf{W}}_{\Phi}$, we provide an alternative initialization approach that considers both weight and activation effects. Specifically, we train each $\mathbf{W}_{\Phi}$ for $200$ iterations to minimize the quantization error of its corresponding linear layer’s output (i.e., $\arg\,\min_{\mathbf{W}_{\Phi}}\|\mathbf{Y}-\hat{\mathbf{Y}}\|^{2}_{F}$). The resulting layer-wise trained $\mathbf{W}_{\Phi}$ is then used to initialize $\Phi$. As shown in Tab. N, both initialization strategies yield similar performance. Therefore, we believe that QVGen is not sensitive to such choices of initialization.
J.3 Complete Results of Tables in Ablation Studies
In Tabs. O to R, we present the complete ablation results across all $8$ dimensions on VBench (Huang et al., 2024b), corresponding to the incomplete versions shown in the ablation study of the main text. These results are consistent with the analyses provided in the main text.
Table O: Complete ablation results of each component. “Naive” denotes naive QAT in a KD-based manner. “-decay” denotes the setting where ${\mathbf{W}}_{\Phi}={\mathbf{L}}{\mathbf{R}}$ is initialized with $r=32$ but not eliminated during QAT. The comparable performance between “-Decay” and $\Phi$” validates that a low-rank setting with $r<d$ (as mentioned in the main text) is sufficient. Furthermore, the negligible performance loss of “++Rank” compared to “-Decay” confirms the effectiveness of our proposed decay strategy.
Method
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Table P: Complete results of different shrinking ratios $\lambda$ for each decay phase. $\lambda=1$ means directly decaying the entire ${\mathbf{W}}_{\Phi}$.
$\quad\quad\lambda\quad\quad$
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Table Q: Complete results of different initial ranks $r$. $r=0$ represents “Naive” in Tab. O.
$\quad\quad r\quad\quad$
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Table R: Complete results of different decay strategies. “Rank” denotes the rank-decay strategy in this work. To be noted, the “Sparse” and “Res. Q.” strategies incur substantially $1.81\times$ and $2.60\times$ GPU days for training compared with the “Rank” approach, respectively (see Sec. F).
Decay
Strategy
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
J.4 Additional Rank-Based Regularization $\bm{\gamma}$}
@@
$\quad\quad{\bm{\gamma}}\quad\quad$
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Appendix K Results for Image Generation
Table T: W4A4 quantization results for SD3-medium (Esser et al., 2024b). We evaluate FID (Heusel et al., 2018) and CLIP score (Hessel et al., 2022) on the MJHQ-3030(Li et al., 2024a) dataset and employ GenEval (Ghosh et al., 2023) to further measure text-image alignment.
@@
Our theoretical insight is broadly applicable to QAT. Additionally, the proposed QAT strategy is independent of model architecture and data type, so it can be transferred to other domains. However, further studies are needed to confirm whether the same strategy and insight can deliver similar significant improvements in other tasks. To be specific, the large gains we report for video generation rely on the observed behavior of the gradient norm and the singular values during QAT for video diffusion. Here, we report the initial results for image generation in Tab. T, which show that our approach can also achieve non-negligible performance enhancement without additional inference overhead. We will explore more about this in the future.
Appendix L Qualitative Results
@@
In this section, we present random samples generated by video DMs without cherry-picking, as exhibited from Figs. A-H. For a detailed comparison, zoom in to closely examine the relevant frames.

### v5

@@ Appendix F Additional Fine-Grained Decay Strategies
In this section, we detail the “Sparse” and “Res. Q.” decay strategies (see Tab. N), which shrink $\Phi$ to $\varnothing$, as follows:
@@
    $\begin{array}[]{ll}{\mathbf{W}}_{\Phi}={\mathcal{Q}}_{4}({\mathbf{W}}_{\Phi})+{\mathcal{Q}}_{4}({\mathbf{E}}_{1})+{\mathcal{Q}}_{4}({\mathbf{E}}_{2})+{\mathcal{Q}}_{4}({\mathbf{E}}_{3}).\end{array}$ (T)
@@
Compared with our rank-decay strategy (with $r=32$), both alternative methods require storing multiple tensors [^9] The “Sparse” strategy additionally requires storing binary masks to enforce sparsity., each with the same shape as the corresponding linear layer’s weight matrix. In particular, the “Res. Q.” strategy incurs substantial memory overhead and necessitates CPU offloading to avoid out-of-memory (OOM) issues. Furthermore, the “Sparse” strategy introduces extra computational cost due to the matrix multiplication between the mask and ${\mathbf{W}}_{\Phi}$. In practice, the “Sparse” and “Res. Q.” strategies consume $20.12$ and $28.78$ GPU days, respectively, whereas our proposed rank-decay requires only $11.11$ GPU days (see Tab. A). More importantly, rank-decay also achieves significantly better performance than both alternatives, as demonstrated in Tab. N.
Appendix G Training Loss Curves
In this section, we present training loss curves across different methods and models. As shown in Fig. A, QVGen and our method in Sec. 3.1 achieve faster and more stable convergence, which supports the effectiveness of the proposed approaches. Note that LSQ (Esser et al., 2020b) uses $\mathbb{E}_{{\bm{x}}_{0},\mathcal{C},\tau}[\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{\tau},\mathcal{C},\tau)\|_{F}^{2}]$ with ${\bm{x}}_{\tau}=\alpha_{\tau}{\bm{x}}_{0}+\sigma_{\tau}{\bm{\epsilon}}$ as its training objective, instead of Eq. (6) used by the remaining distillation-based methods.
(a) CogVideoX-22(Yang et al., 2025)
(b) Wan 1.31.3(WanTeam et al., 2025)
Figure A: Training loss vs. #steps across different video DMs and $4$-bit QAT methods. $\Phi$” denotes our approach in Sec. 3.1
Appendix H Further Analyses of Gradient Norm in Video Generation QAT
H.1 Video Generation QAT vs. Image Generation QAT
First, we conduct experiments to compare the gradient norm between image generation QAT and video generation QAT. In Tab. C, we employ W4A4 Q-DM (Li et al., 2023b) to quantize the diffusion models under the same settings. We observe that reducing the gradient norm during QAT, ignored by previous research (Li et al., 2023b; He et al., 2024; Esser et al., 2020b), is far more critical for video generation QAT than for image generation QAT. As Tab. C shows, with a similar parameter count and the same QAT recipe, video diffusion reaches a significantly larger gradient norm than image diffusion. This leads to much more unstable optimization (Xie et al., 2024) in training. We believe the phenomenon happens because video generation introduces complicated temporal modeling, which eventually makes quantization for video diffusion more challenging than image diffusion.
Table C: Average gradient norm comparison. SD3-medium (Esser et al., 2024a) is an advanced diffusion model for image generation. We train SD3-medium for $2K$ steps with $16K$ images from the LAION-55B dataset (Schuhmann et al., 2022) on 88 $H100$ GPUs. “Avg. $\|{\bm{g}}_{t}\|_{2}$” is the average of the gradient norm across training steps.
@@ Appendix H Further Analyses of Gradient Norm in Video Generation QAT
H.2 Impact of Motion Dynamics on Gradient Norm
Here, we study how the motion dynamics of video generation affect the gradient norm. Specifically, using UniMatch (Xu et al., 2023a), we compute an optical‑flow score as a motion difference score for each clip and split the training data (Sec. 4.1) into high‑motion and low‑motion subsets, each with $8K$ videos. In Tab. D, the model trained on the low‑motion subset shows a much lower average gradient norm and a better Imaging Quality score, but its Dynamic Degree is lower than that of the model
Table D: Results between different training videos. We employ W4A4 QVGen with the same configurations as those in Tab. 1 to quantize CogVideoX-22(Yang et al., 2025).
@@ H.2 Impact of Motion Dynamics on Gradient Norm
trained on the high‑motion subset. This pattern suggests that a high-motion training set makes QAT harder and less stable (i.e., higher gradient norm), lowering static quality but boosting motion quality. The reverse holds for low‑motion clips. To be noted, with a mixed dataset (containing low-motion ++ high-motion) in the $4$-th row of Tab. D, the quantized model generalizes well in producing either high‑motion or low‑motion content. Therefore, although a low-motion dataset in training can cause slightly lower $\|{\bm{g}}_{t}\|_{2}$ (i.e., better training convergence), it is necessary to properly add high-motion data to improve motion quality.
H.3 Directly Regulate the Gradient Norm
In this subsection, we study the effect of directly constraining the gradient norm during QAT. We use torch.nn.utils.clip_grad_norm_ to rescale the gradients. As shown in Tab. E, reducing the clipping threshold from $1.0$ to $0.5$ improves performance, which supports the benefit of controlling gradient norms for video generation QAT. However, a threshold of $0.1$ causes clear performance degradation, likely because the quantized model is updated too weakly or aggressive gradient clipping disrupts normal QAT. This highlights the need for more principled ways to reduce the gradient norm.
Table E: W4A4 results for Wan 1.31.3B under different thresholds for gradient clipping. “1.0” corresponds to our baseline Q-DM (Li et al., 2023b). We use a clipping threshold of $1.0$ for all other experiments in the paper.
Grad. Clipping Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
1.0 60.40 52.50 76.67 13.28 21.63
0.5 60.58 52.95 76.50 13.28 21.71
0.1 56.68 51.06 71.00 12.78 21.42
Appendix I Combination with SVDQuant
In Tab. F, we show that our method can be combined with the current SOTA PTQ method SVDQuant (Li et al., 2025). We first apply SVDQuant to obtain a weight-modified DM, the quantization parameters, and the low-rank matrices, which we reuse as $\Phi$. We then run QVGen, which progressively removes $\Phi$. This combination yields further gains, likely due to the strong initialization from SVDQuant. We also evaluate an option that updates only the quantization parameters and $\Phi$ within this combination. It achieves sizable improvements over SVDQuant alone, but it still falls short of QVGen when model weights are updated. This suggests that QAT, which trains model weights, is currently important for video generation quantization. We will continue to explore more efficient ways to quantize video DMs while preserving performance.
Table F: W4A4 results of the combination with SVDQuant (Li et al., 2025) for Wan 1.3B (WanTeam et al., 2025). $\heartsuit$” denotes we freeze the weights of the DM and only finetune the quantization parameters and the introduced $\Phi$. $\clubsuit$” means we employ a more fine-grained and performance-friendly quantization setting as in SVDQuant’s paper (details can be found in Sec. D).
Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
Full Prec. 64.30 58.21 70.28 28.05 24.67
SVDQuant (Li et al., 2025) 57.57 46.30 72.22 12.73 21.91
QVGen 63.08 54.67 77.78 15.32 23.01
QVGen w/ SVDQuant (Li et al., 2025) 63.64 56.23 77.42 17.65 23.89
QVGen w/ SVDQuant (Li et al., 2025) 61.38 52.76 75.85 14.12 22.47
Appendix J Comparison with baselines on Additional Metrics
Table G: Additional W4A4 performance comparison across different quantization methods. We employ the same models as those in Tab. 1.
@@ Appendix I Combination with SVDQuant
We also evaluate the similarity between videos generated by different quantization methods and those generated by BF16 models on VBench (Huang et al., 2024c) captions. Specifically, we employ PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity) (Wang et al., 2004), and LPIPS (Learned Perceptual Image Patch Similarity) (Zhang et al., 2018). In Tab. G, QVGen substantially outperforms the baselines on these metrics.
Appendix K Comparison with Baselines for Huge DMs
Table H: Performance comparison for huge DMs across different methods on VBench (Huang et al., 2024c). We employ W4A4 CogVideo-X 55(Yang et al., 2025) here.
Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
We include a comparison on VBench for large-scale CogVideoX‑1.5 55B in Tab. H. QVGen again surpasses all baselines, just as it does on smaller models, underscoring its strength across a wide range of model sizes. Limited resources currently prevent us from adding more baselines for these large-scale models, but we plan to do so in future work.
Appendix L More Ablation Studies
L.1 Robustness of Rank-Decay Schedule across Different Annealing Functions
In this section, we test $5$ different annealing functions for $u$. The results in Tab. I reveal that all the functions yield comparable performance, highlighting the robustness of our approach.
Table I: Results of different annealing factors $u$. All of them decay from $1$ to 0. We employ “Cosine” in this work.
$\quad u\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
L.2 Initialization Methods for Auxiliary Modules $\Phi$
Besides employing ${\mathbf{W}}-{\mathcal{Q}}_{b}({\mathbf{W}})$ to initialize ${\mathbf{W}}_{\Phi}$, we provide an alternative initialization approach that considers both weight and activation effects. Specifically, we train each $\mathbf{W}_{\Phi}$ for $200$ iterations to minimize the quantization error of its corresponding linear layer’s output (i.e., $\arg\,\min_{\mathbf{W}_{\Phi}}\|\mathbf{Y}-\hat{\mathbf{Y}}\|^{2}_{F}$). The resulting layer-wise trained $\mathbf{W}_{\Phi}$ is then used to initialize $\Phi$. As shown in Tab. J, both initialization strategies yield similar performance. Therefore, we believe that QVGen is not sensitive to such choices of initialization. Moreover, we also consider two deliberately suboptimal initialization approaches: zero initialization (i.e., $\mathbf{0}$”) and initialization with parameters randomly sampled from a normal distribution (i.e., “Random”). The $\mathbf{0}$” scheme leads to a slight performance degradation, while “Random” causes a more noticeable performance drop. We attribute this to the fact that $\mathbf{0}$” does not compensate for quantization errors, whereas “Random” further injects noise into the model.
Table J: W4A4 results of different initialization strategies for ${\mathbf{W}}_{\Phi}$.
Init. Strategy Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
$\mathbf{0}$ 62.80 54.59 77.69 15.28 22.98
Random 54.41 44.14 34.44 3.13 10.17
L.3 Complete Results of Tables in Ablation Studies
In Tabs. K to N, we present the complete ablation results across all $8$ dimensions on VBench (Huang et al., 2024c), corresponding to the incomplete versions shown in the ablation study of the main text. These results are consistent with the analyses provided in the main text.
Table K: Complete ablation results of each component. “Naive” denotes naive QAT in a KD-based manner. “-decay” denotes the setting where ${\mathbf{W}}_{\Phi}={\mathbf{L}}{\mathbf{R}}$ is initialized with $r=32$ but not eliminated during QAT. The comparable performance between “-Decay” and $\Phi$” validates that a low-rank setting with $r<d$ (as mentioned in the main text) is sufficient. Furthermore, the negligible performance loss of “++Rank” compared to “-Decay” confirms the effectiveness of our proposed decay strategy.
Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
Table L: Complete results of different shrinking ratios $\lambda$ for each decay phase. $\lambda=1$ means directly decaying the entire ${\mathbf{W}}_{\Phi}$.
$\quad\quad\lambda\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
Table M: Complete results of different initial ranks $r$. $r=0$ represents “Naive” in Tab. K.
$\quad\quad r\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
Table N: Complete results of different decay strategies. “Rank” denotes the rank-decay strategy in this work. To be noted, the “Sparse” and “Res. Q.” strategies incur substantially $1.81\times$ and $2.60\times$ GPU days for training compared with the “Rank” approach, respectively (see Sec. F).
Decay Strategy Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
L.4 Additional Rank-Based Regularization $\bm{\gamma}$}
@@
$\quad\quad{\bm{\gamma}}\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
L.5 Duration of Each Decay for $\bm{\gamma}$}
Here, we discuss the situation if $\Phi$’s rank is diminished too rapidly relative to the schedule $\mathbf{\gamma}$. In this case, we believe that keeping the redundant parts (that is, the rank-diminished components in $\Phi$) during QAT does not harm performance and can even lead to a small improvement. This is mainly because the slow change of $u$ in $\mathbf{\gamma}$ ($1\rightarrow 0$) increases the training time by making each decay phase longer. The results in the following table support this intuition.
Table P: W4A4 quantization results across different durations of each decay phase for Wan 1.31.3B. We control the duration to determine the changing speed of the schedule $\mathbf{\gamma}$. When applying a long duration, we suggest $\Phi$’s rank is diminished (as an intrinsic behavior) rapidly relative to the schedule $\mathbf{\gamma}$.
Duration (Epoch) Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
$3/4$ 63.08 54.67 77.78 15.32 23.01
11 63.07 55.09 77.54 15.29 23.03
$3/2$ 63.11 54.58 78.15 15.36 23.01
L.6 Weight-Only Quantization vs. Activation-Only Quantization
As demonstrated in Fig. B, activation-only quantization causes severe degradation in video generation quality compared with weight-only quantization under the $4$-bit setting. This indicates that the activations of video generation models are much harder to quantize than the weights. Similar observations have also been reported in previous studies (Zhao et al., 2025a; Tian et al., 2024).
(a) BF16
(b) BF16
(c) W4A16
(d) W4A16
(e) W16A4
(f) W16A4
Figure B: Performacne for Weight-only quantization vs. activation-only quantization. We employ Min-max (Nagel et al., 2021) per-channel weight quantization and per-token activation quantization for (Left) CogVideoX-22(Yang et al., 2025) and (Right) Wan 1.31.3(WanTeam et al., 2025).
Appendix M Results for Image Generation
Our theoretical insight is broadly applicable to QAT. Additionally, the proposed QAT strategy is independent of model architecture and data type, so it can be transferred to other domains. However, further studies are needed to confirm whether the same strategy and insight can deliver similar significant improvements in other tasks. To be specific, the large gains we report for video generation rely on the observed behavior of the gradient norm and the singular values during QAT for video diffusion. Here, we report the initial results for image generation in Tab. Q, which show that our approach can also achieve non-negligible performance enhancement without additional inference overhead. We will explore more about this in the future.
Table Q: W4A4 quantization results for SD3-medium (Esser et al., 2024b). We evaluate FID (Heusel et al., 2018) and CLIP score (Hessel et al., 2022) on the MJHQ-3030(Li et al., 2024a) dataset and employ GenEval (Ghosh et al., 2023) to further measure text-image alignment.
@@
Appendix N Profiling and Projected Gains from Kernel Fusion
Table R: Latency breakdown (ms) and INT4 GEMM throughput on $A800$. For ease of analysis, we adopt the shapes from Wan 1.31.3B. A DiT block for the model is composed of Self-Attention, Cross-Attention, and FFN. “q1/k1/v1/o1” and “q2/k2/v2/o2” denote projections in Self-Attention and Cross-Attention, respectively.
Op. $(M,N,K)$ Quant INT4GEMM DeQuant Total Quant (%) INT4GEMM (%) DeQuant (%) TOPS
q1/k1/v1/o1/q2/o2 $(32760,1536,1536)$ 0.041 0.186 0.114 0.341 12.0 54.5 33.5 831.08
k2/v2 $(512,1536,1536)$ 0.001 0.007 0.002 0.010 11.2 72.3 16.5 345.13
up_proj $(32760,8960,1536)$ 0.041 1.246 0.661 1.948 2.10 64.0 33.9 721.76
down_proj $(32760,1536,8960)$ 0.232 1.031 0.117 1.380 16.8 74.7 8.48 872.27
To better understand the efficiency bottlenecks in our current non-fused INT4 implementation, we profile representative transformer operators on an NVIDIA $A800$ (SM8080) by annotating stages such as activation quantization (Quant), INT4 general matrix multiplication (INT4GEMM), and dequantization (DeQuant) with NVTX ranges (torch.cuda.nvtx.range_push/pop). For each operator in Tab. R, we report its GEMM dimensions $(M,K,N)$ (i.e., $M,K]\!\times\![K,N]$), per-stage latency, the fraction of the total operator time, and the effective INT4 GEMM throughput computed as
$\begin{array}[]{cc}\text{TOPS}=\frac{2MNK}{t_{\texttt{INT4GEMM}}}\,\div\,10^{12},\end{array}$ (U)
where $t_{\texttt{INT4GEMM}}$ is the measured INT4GEMM time in seconds. Across all tested shapes, the INT4 GEMM kernels achieve $345\text{--}872$ TOPS, within the same order of magnitude as the $A800$’s INT4 tensor-core peak ($1248$ TOPS). However, the surrounding non-GEMM stages (activation quantization and dequantization) still account for $25\text{--}45\%$ of the operator latency, primarily due to extra global memory traffic and the absence of fused epilogues. Given the measured GEMM time fraction $G$ and assuming kernel fusion removes a fraction $r$ of non-GEMM overhead (e.g., fusing quant/dequant and avoiding intermediate reads/writes), the achievable speedup is approximated by
$\begin{array}[]{ll}\rho\approx\frac{1}{G+(1-G)(1-r)}.\end{array}$ (V)
Using our NVTX-derived $G$ values, $r=0.6$ yields $\rho\!\approx\!1.18{\times}$ (down_proj), $1.20{\times}$ (k2/v2), $1.28{\times}$ (up_proj), $1.38{\times}$ (q1/k1/v1/o1/q2/o2); and $r\approx 0.8$ yields $\rho\!\approx\!1.25{\times}$, $1.28{\times}$, $1.40{\times}$, and $1.57{\times}$, respectively. These results indicate that a $1.2\text{--}1.6{\times}$ per-layer speedup is a realistic target once fusion is introduced.
Appendix O Breakdown Latency Analysis
Because a video DiT is implemented as a stack of identical blocks, we report the latency breakdown of a single DiT block to estimate its end-to-end impact (see Tabs. U-U). Within this block, attention computation ($51.8\%$) is the dominant cost, while linear projections ($24.5\%$) account for a large share of the remaining latency. To be noted, components other than linear projections can be accelerated by orthogonal strategies:
  • Attention: Sparse attention, such as SVG (Xi et al., 2025), achieves a $1.73\times$ speedup for Self-Attention computation ($31.48$ vs. $18.23$).
  • Other: This category is largely composed of memory-bound operations, including RoPE, norm, and reshape, etc. These operations often launch many small kernels, so techniques such as CUDA Graphs and torch.compile can reduce dispatch overhead and enable more effective kernel fusion. Additionally, combined with fused and layout-aware kernels (Xi et al., 2025), the runtime of this category can be reduced by $5.59\times$ ($14.64$ vs. $2.620$).
With these strategies applied, linear projections occupy a non-trivial $41.38\%$ of the block runtime. Therefore, reducing the latency of the linear projections is an important step toward further end-to-end speedups. In this work, W4A4 quantization achieves a $2.52\times$ speedup for these linear projections ($15.14$ vs. $5.991$). In addition, we plan to extend QVGen to W4A4 attention quantization, which can further accelerate attention computation.
Table S: Latency breakdown (ms) for a DiT block (implemented in torch) on $A800$ (Wan 1.31.3B).
Component Time (ms) Share (%)
Attention 32.08 51.8
Linear Projections 15.14 24.5
Other 14.64 23.7
Table T: Latency breakdown (ms) for linear projections.
Linear projections Time (ms)
q1/k1/v1/o1/q2/o2 0.977
k2/v2 0.038
up_proj 5.283
down_proj 3.913
Table U: Latency breakdown (ms) for attention.
Attention Time (ms)
Self-Attention 31.48
Cross-Attention 0.597
Appendix P Qualitative Results
@@
In this section, we present random samples generated by video DMs without cherry-picking, as exhibited from Figs. C-J. For a detailed comparison, zoom in to closely examine the relevant frames.